OpenStudy (tiffany_rhodes):

Use Euler's formula to write the expression e^(2+7i) in the form: a+bi. Round a and b to four decimal places. I cannot get the answer for the life of me!

OpenStudy (anonymous):

\[e^{2+7i}=e^2\left(\cos7+i\sin 7\right)\] Use a calculator to find approximate values of \(a=e^2\cos7\) and \(b=e^2\sin7\). Remember to use radians, not degrees.